OpenStudy (anonymous):

Oops, accidentally closed instead of going to/bumping Consider the function f(x) = 8/x^3 - 4/x^5. Let F(x) be the antiderivative of f(x) with F(1) = 0. Then F(5) equals

OpenStudy (anonymous):

Have to go but if anyone can help with this I will be checking in when I can

OpenStudy (anonymous):

\[f(x)={8\over x^3}-{4\over x^5}=8x^{-3}-4x^{-5}\\ F(x)=\int(8x^{-3}-4x^{-5})dx=8{x^{-2}\over-2}-4{x^{-4}\over-4}+C\\ F(x)=-4x^{-2}+x^{-4}+C\\ F(1)=0=-4+1+C\implies C=3\\ F(5)=-4(5^{-2})+(5^{-4})+3\\ \implies F(5)={-4\over25}+ {1\over625}+3=? \]